If the reserve ratio decreased from 20 percent to 10 percent, which of the following would happen to the money multiplier?
a.
It would rise from 10 to 20.
b.
It would rise from 5 to 10.
c.
It would fall from 10 to 5.
d.
It would fall from 20 to 5.
13. Which statement best describes the outcome of a decrease in the bank rate?
a.
Banks will borrow less from Bank of Canada, so reserves increase.
b.
Banks will borrow less from Bank of Canada, so reserves decrease.
c.
Banks will borrow more from Bank of Canada, so reserves increase.
d.
Banks will borrow more from Bank of Canada, so reserves decrease.
14. Which statement best describes the process of open-market sales conducted by the Bank of Canada?
a.
The Bank of Canada sells Treasury bills, which increases the money supply.
b.
The Bank of Canada sells Treasury bills, which decreases the money supply.
c.
The Bank of Canada borrows from member banks, which increases the money supply.
d.
The Bank of Canada lends money to member banks, which decreases the money supply.
15. What is a characteristic of Scotiabank?
a.
It can issue currency.
b.
It is part of the “big 5” commercial banks group.
  
c.
It acts as a central bank for Nova Scotia.
d.
It is owned by the Canadian government.
16. If a central bank wanted to increase the money supply, what would it most likely do?
a.
It would make open-market purchases and lower the bank rate.
b.
It would make open-market sales and lower the bank rate.
c.
It would make open-market purchases and raise the bank rate.
d.
It would make open-market sales and raise the bank rate.
17. Which statement best defines the bank rate?
a.
It is the interest rate the Bank of Canada charges banks.
b.
It is one divided by the difference between one and the reserve ratio.
c.
It is the interest rate banks receive on reserve deposits with Bank of Canada.
d.
It is the interest rate that banks charge on overnight loans to other banks.
Answer #1

Due to presence of HOMEWORKLIB POLICY, I am answering four questions.

12.

Ans: b.

Explanation: Initial multiplier = 1 / 0.2 = 5

New multiplier = 1 / 0.1 = 10

13.

Ans: c.

Explanation: A bank rate is the interest rate at which a nation's central bank lends money to domestic banks, often in the form of very short-term loans.

14.

Ans: b.

Explanation: Selling treasury bills imply taking money out from economy.

16.

Ans: a

Explanation: Purchasing operation implies injecting money into the economy and bank rate is the interest rate at which a nation's central bank lends money to domestic banks, often in the form of very short-term loans.

If you are satisfied with the answer, please provide a positive rating. Feel free to comment in case of queries.

Have a nice day ahead!